# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 5 Hierarchy of skills
Regardless of whether the prompting input to LLMs is in natural language or PDDL format, the hierarchy of skills possessed by LLMs still exhibits instability (Silver et al. (2022)). Hence, researchers are exploring approaches that integrate LLMs with classical PDDL-based planning methods for the hierarchy of skills. LLM+P (Liu et al. (2023a)) utilize LLMs to translate natural language into PDDL and input into a classical planner for the hierarchy of skills. Xie et al. (2023b) indicate that LLMs exhibit greater efficacy in translation tasks as opposed to planning.
@@ 6 State
3D reconstruction involves capturing both the shape and appearance of an object or scene (Wikipedia Contributors (2025)). 3D reconstruction methods are divided into passive and active types (Butime et al. (2006)). Active methods involve contact or project some form of energy onto the object, such as Laser Scanning (Butime et al. (2006)), X-ray (Maken and Gupta (2023)). These devices have high accuracy, but they are usually expensive. Therefore, various studies focus on 3D reconstruction using consumer RGBD cameras, such as Microsoft Kinect, Intel RealSense, Google Tango, and ORBBEC Gemini (Li et al. (2022)).
@@
These consumer cameras typically use principles such as structured light, time of flight, and traditional photometric stereo for depth estimation (Zhou et al. (2024a)) and 3D representation can be generated by registering them using camera poses (Huang et al. (2024b)). However, when surfaces are shiny, bright, transparent, textureless or distant from the camera, depth images produced by consumer cameras are often noisy and incomplete. Several studies have addressed this challenging problem by learning to restore depth images (Dai et al. (2022); Fang et al. (2022); Sajjan et al. (2020)). However, the correct depth information may already be lost in the original depth. ASGrasp (Shi et al. (2024)) demonstrates that 3D reconstruction using raw multi-view images from consumer cameras is better than restoring the original depth. Many studies use a single image for 3D reconstruction (Fu et al. (2021)). However, a single image loses a significant amount of information, resulting in lower accuracy. Despite this, the zero-shot capability of current single image 3D reconstruction has led to its widespread application in simulation scene generation (in Sec. 8.3).
@@
The 3D representation for 3D reconstruction can be expressed as explicit and implicit expressions (Zhou et al. (2024a)). Explicit expressions include point clouds (Shi et al. (2024)), voxels (Jiang et al. (2021)), and meshes (Wen et al. (2019)). The three representations can be converted into each other (Jiang et al. (2021)), but each has its own advantages. A point cloud is made up of discrete points in space, providing flexibility in processing. A voxel can store spatial information inside an object but comes with high space complexity. A mesh uses triangle meshes to represent complex shapes and details accurately, such as deformation (Wen et al. (2019)). It ensures the projection is always convex, making it easier to rasterize (Zhou et al. (2024a)). Implicit expressions represent 3D geometry using a function, such as signed distance function (SDF) (Chabra et al. (2020)), occupancy field (Jiang et al. (2021)), and radiance field (Mildenhall et al. (2021)). They offer differentiability and efficient storage, making them a powerful tool. In contrast, explicit expressions tend to be more intuitive.
@@
GIGA (Jiang et al. (2021)) points out that manipulation requires a fine-grained understanding of local geometry details. Implicit representations, due to their continuous and differentiable nature, can represent smooth surfaces at high resolution. As a result, there is a growing research using implicit representations for manipulation tasks (Dai et al. (2023); Lu et al. (2024)). Current state-of-the-art methods for representing scenes using implicit representations are mainly divided into Nerf-based (Wang et al. (2023b)) and 3DGS-based approaches (Kulhanek and Sattler (2024)). Compared to Nerf (Mildenhall et al. (2021)), 3DGS offers better real-time performance (Kerbl et al. (2023)). However, these implicit 3D representations currently lack scene semantics and not easily editable for 3D modifications (Bai et al. (2024)).
@@
As for scene semantics, semantic-NeRF (Zhi et al. (2021)) employs manually annotated semantic labels to jointly encode semantics, appearance, and geometry using NeRF. Manual annotation is time-consuming and labor-intensive. Due to the foundation model’s robust open-set capability for objects, DFF (Shen et al. (2023)), CLIP-Fields (Shafiullah et al. (2022)) and LERF (Kerr et al. (2023)) employ CLIP image encoder to extract features from multi-view 2D images for NeRF (Mildenhall et al. (2021)) reconstruction. These features are integrated as part of the output of the NeRF network, enriching the semantic information of the reconstructed 3D scenes. When a text prompt is provided, the features output by the CLIP text encoder can be compared with the CLIP image features output by NeRF to form a relevancy map. This relevancy map can support downstream tasks, such as semantic scene completion and object localization (Ha and Song (2022)). Since CLIP can only provide image-level features, the relevancy map lacks precise pixel-level object boundary information. 3DOVS (Liu et al. (2023c)) incorporates DINO features into the NeRF output to distill object boundary information. OV-NeRF (Liao et al. (2024)) addresses the issues of coarse relevancy maps and view-inconsistent relevancy maps through SAM and cross-view self-enhancement. FMGS (Zuo et al. (2025)) transfers this concept from NeRF to 3DGS, achieving 851X faster inference. Although foundation models, such as CLIP and DINO, enable 3D open-set semantic scene understanding, the performance is limited by the foundation models themselves. For example, CLIP is constrained by the bag-of-words limitation (Kerr et al. (2023)).
@@
The image features output by NeRF can be used to build a relevancy map. They can also be lifted into 3D space through multi-view images, serving as 3D features for downstream tasks (Ze et al. (2023)). 3D-LLM (Hong et al. (2023b)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023b)), endowing 3D features with semantic information.
@@
For implicit 3D editing, some methods use human scribbles to edit 3D shape and appearance (Zhang et al. (2023d); Schwarz et al. (2020); Li and Pan (2024); Liu et al. (2021)). However, this approach is not intuitive enough. With the development of foundation models, many methods for implicit 3D editing using text prompts have emerged. CLIP-NeRF (Wang et al. (2022a)) integrates semantic features extracted by CLIP into NeRF reconstruction to change object shape and appearance during rendering. However, CLIP-based approaches cannot precisely modify specific local regions. Instruct-NeRF2NeRF (Haque et al. (2023)) utilizes InstructPix2Pix (Brooks et al. (2023)) to iteratively edit multiview input images and optimize the underlying scene in NeRF. This process produces a refined 3D scene that adheres to the edit instruction. However, InstructPix2Pix modifies the entire image. As a result, regions that are not desired may also be altered. DreamEditor (Zhuang et al. (2023)) uses Dreambooth (Ruiz et al. (2023)) to generate 2D editing masks. These masks are then converted into 3D editing regions through back projection. This approach enables precise local editing.
@@ 6.2 Pose Estimation
Object pose estimation can be divided into marker-based and markerless methods (Karashchuk et al. (2021)). Marker-based methods require attaching passive or active markers (Cassinis and Tampalini (2007)) to the object. These methods achieve high accuracy in pose estimation. For example, the NDI Polaris Vega XT, commonly used in medical robotics, can achieve an accuracy of 0.12 mm RMS (NDI (2024)). However, in unstructured environments, it is not feasible to attach specific markers to every object. Therefore, achieving object pose estimation in unstructured markerless environment is necessary.
@@
From the perspective of generalization, pose estimation methods can be classified into instance-level, category-level, and unseen object approaches (Liu et al. (2024c)). Instance-level methods can estimate pose accurately for specific object instances on which they are trained. However, they struggle with novel objects. To improve the model’s adaptation for pose estimation of novel objects, category-level approaches use geometric priors from objects of the same category to estimate the pose of the novel object without requiring its 3D model (Wang et al. (2019)). Unseen object approaches typically rely on the 3D model of the novel object to estimate its pose (Caraffa et al. (2024)).
@@
Category-level and unseen object approaches can also be primarily classified into model-free and model-based approaches (Liu et al. (2024c)). Model-free methods do not require prior knowledge of the object’s 3D model. These methods typically regress the object pose using neural network (Guan et al. (2024)). However, these methods require large amounts of data for the neural network to learn the geometric priors of the object. In contrast, model-based methods use a known 3D model of the object and usually lead the BOP benchmark for object pose estimation (Burde et al. (2024)). However, obtaining accurate object 3D models in the real world is not easy. The advancement of multi-view image 3D reconstruction technology bridges the gap between model-based and model-free real-world applications (Burde et al. (2024)).
@@
The input modalities for the model-based approach include RGB, depth, and RGBD. Currently, the RGBD modality leads the BOP benchmark for object pose estimation. The optimization goals are primarily divided into three parts: 2D-2D correspondences followed by regression (Nguyen et al. (2024)), 2D-3D correspondences followed by PnP (Li et al. (2023b); Ausserlechner et al. (2024)), and 3D-3D correspondences followed by least squares fitting (Lin et al. (2024b); Caraffa et al. (2024)). However, pose estimation accuracy remains a challenge when dealing with occlusion, specularity, symmetry and textureless objects (Guan et al. (2024)). Many methods use the predicted pose as a coarse result and refine it to obtain a fine result (Labbé et al. (2022); Wen et al. (2024a); Moon et al. (2024)).
@@
Figure 6: Foundation Models for State. The foundation models have three main applications in 3D reconstruction: 3D open-set semantic scene understanding, lifting 2D features to 3D space, and implicit 3D editing. In 3D open-set semantic scene understanding, the main pipeline is to use image features extracted by the VFM encoder and VLM image encoder as input for NeRF. Then, semantic text features extracted by the VLM language encoder are used in conjunction with the image features from NeRF to generate a relevancy map through a relevancy extractor (Kerr et al. (2023)). This relevancy map can support downstream tasks, such as semantic scene completion and object localization (Ha and Song (2022)). As for lifting, using the VLM image encoder to extract features from 3D data multi-view images and lift them into 3D features can incorporate semantic information into the 3D features. The lifting methods include direct reconstruction, gradSLAM, and Neural Field (Hong et al. (2023b)). For implicit 3D editing, the current mainstream pipeline is to input the image rendered by NeRF and the editing prompt into the VGM to generate the updated image. The updated image is then fed back into NeRF for training, modifying NeRF’s radiance field representation of the 3D scene (Haque et al. (2023)). Pose estimation with foundation models achieves state-of-the-art results (Caraffa et al. (2024)). The main method is 2D-lifting-3D. It extracts texture features from the object model and observation RGBD image. LMM extracts geometric features from the object model and observation RGBD image. The fused features are then used to estimate the 6D pose through 3D-3D correspondences.
@@
The pose estimation of moving objects mainly involves two methods. 1) Some single-image 6D pose estimation methods are fast and re-estimate poses from scratch for each frame. However, this approach is inefficient and results in less coherent estimations (Wen and Bekris (2021)). 2) Pose tracking utilizes temporal cues to improve pose prediction. It enhances efficiency, smoothness, and accuracy in video sequences. Current pose-tracking methods are mainly divided into probabilistic tracking (Deng et al. (2021); Issac et al. (2016); Stoiber et al. (2022)) and optimization-based tracking (Li et al. (2018); Wen et al. (2020); Lin et al. (2022); Wang et al. (2020a)). Pose tracking faces challenges mainly from motion blur, incremental error drift, and occlusion. To address these issues, BundleSDF (Wen et al. (2023a)) and BundleTrack (Wen and Bekris (2021)) use an online pose graph optimization process.
@@
There are some research integrating foundation models into pose estimation. As for category-level, OV9D (Cai et al. (2024)) utilizes DINO and VQVAE to extract visual features from images, while CLIP is used to extract text features from category prompts. These features are then fed into the Stable Diffusion UNet (Rombach et al. (2022)) to generate a normalized object coordinate space (NOCS) map. This method achieves generalizability to unseen categories and enables open-set pose estimation. In unseen object pose estimation with foundation models, FoundationPose (Wen et al. (2024a)) utilizes LLM-aided synthetic data generation at scale to ensure strong generalizability for novel object pose estimation & tracking. SAM-6D (Lin et al. (2024b)) and ZS6D (Ausserlechner et al. (2024)) leverage SAM to generate valid proposals, enabling zero-shot 6D pose estimation. FreeZe (Caraffa et al. (2024)) employs frozen GeDi (Poiesi and Boscaini (2022)) and DINO (Caron et al. (2021)) to extract both geometric and visual features from the query object model and the target object’s RGBD observation image. It then uses 3D-3D fused feature correspondences to obtain the 6D pose. Due to the foundation models’ robust capability in discriminative feature extraction, FreeZe achieves state-of-the-art results without the need for any data or training. Overall, foundation models primarily improve generalization for novel object pose estimation in three areas: data , object recognition, and feature extraction. However, the performance is limited by the foundation models themselves. For example, foundation models are large in size (Caraffa et al. (2024)) and SAM may hallucinate in object segmentation (Kirillov et al. (2023)).
@@
Following Fig. 6, VLM and VFM assist implicit 3D reconstruction by generating relevancy maps that include semantic information (Kerr et al. (2023)). They can also be employed in 2D-to-3D lifting to extract 3D features, encompassing texture, semantic, and spatial information (Hong et al. (2023b)). VGM aids in generating edited images and modifying 3D scenes based on these edited images (Haque et al. (2023)). FreeZe (Caraffa et al. (2024)) achieves state-of-the-art result in pose estimation by extracting discriminative features through 2D-to-3D lifting and LMM.

### v7

@@ 5 Hierarchy of skills
Regardless of whether the prompting input to LLMs is in natural language or PDDL format, the hierarchy of skills possessed by LLMs still exhibits instability (silver2022pddl). Hence, researchers are exploring approaches that integrate LLMs with classical PDDL-based planning methods for the hierarchy of skills. LLM+P (liu2023llm) utilize LLMs to translate natural language into PDDL and input into a classical planner for the hierarchy of skills. xie2023translating indicate that LLMs exhibit greater efficacy in translation tasks as opposed to planning.
@@ 6 State
3D reconstruction involves capturing both the shape and appearance of an object or scene (wiki2025). 3D reconstruction methods are divided into passive and active types (butime20063d). Active methods involve contact or project some form of energy onto the object, such as Laser Scanning (butime20063d), X-ray (maken20232d). These devices have high accuracy, but they are usually expensive. Therefore, various studies focus on 3D reconstruction using consumer RGBD cameras, such as Microsoft Kinect, Intel RealSense, Google Tango, and ORBBEC Gemini (li2022high).
@@
These consumer cameras typically use principles such as structured light, time of flight, and traditional photometric stereo for depth estimation (zhou2024comprehensive) and 3D representation can be generated by registering them using camera poses (huang2024surface). However, when surfaces are shiny, bright, transparent, textureless or distant from the camera, depth images produced by consumer cameras are often noisy and incomplete. Several studies have addressed this challenging problem by learning to restore depth images (dai2022domain; fang2022transcg; sajjan2020clear). However, the correct depth information may already be lost in the original depth. ASGrasp (shi2024asgrasp) demonstrates that 3D reconstruction using raw multi-view images from consumer cameras is better than restoring the original depth. Many studies use a single image for 3D reconstruction (fu2021single). However, a single image loses a significant amount of information, resulting in lower accuracy. Despite this, the zero-shot capability of current single image 3D reconstruction has led to its widespread application in simulation scene generation (in Sec. 8.3).
@@
The 3D representation for 3D reconstruction can be expressed as explicit and implicit expressions (zhou2024comprehensive). Explicit expressions include point clouds (shi2024asgrasp), voxels (jiang2021synergies), and meshes (wen2019pixel2mesh). The three representations can be converted into each other (jiang2021synergies), but each has its own advantages. A point cloud is made up of discrete points in space, providing flexibility in processing. A voxel can store spatial information inside an object but comes with high space complexity. A mesh uses triangle meshes to represent complex shapes and details accurately, such as deformation (wen2019pixel2mesh). It ensures the projection is always convex, making it easier to rasterize (zhou2024comprehensive). Implicit expressions represent 3D geometry using a function, such as signed distance function (SDF) (chabra2020deep), occupancy field (jiang2021synergies), and radiance field (mildenhall2021nerf). They offer differentiability and efficient storage, making them a powerful tool. In contrast, explicit expressions tend to be more intuitive.
@@
GIGA (jiang2021synergies) points out that manipulation requires a fine-grained understanding of local geometry details. Implicit representations, due to their continuous and differentiable nature, can represent smooth surfaces at high resolution. As a result, there is a growing research using implicit representations for manipulation tasks (dai2023graspnerf; lu2024manigaussian). Current state-of-the-art methods for representing scenes using implicit representations are mainly divided into Nerf-based (wang2023masked) and 3DGS-based approaches (kulhanek2024nerfbaselines). Compared to Nerf (mildenhall2021nerf), 3DGS offers better real-time performance (kerbl20233dgaussiansplattingrealtime). However, these implicit 3D representations currently lack scene semantics and not easily editable for 3D modifications (bai2024survey).
@@
As for scene semantics, semantic-NeRF (zhi2021place) employs manually annotated semantic labels to jointly encode semantics, appearance, and geometry using NeRF. Manual annotation is time-consuming and labor-intensive. Due to the foundation model’s robust open-set capability for objects, DFF (shen2023distilled), CLIP-Fields (shafiullah2022clip) and LERF (kerr2023lerf) employ CLIP image encoder to extract features from multi-view 2D images for NeRF (mildenhall2021nerf) reconstruction. These features are integrated as part of the output of the NeRF network, enriching the semantic information of the reconstructed 3D scenes. When a text prompt is provided, the features output by the CLIP text encoder can be compared with the CLIP image features output by NeRF to form a relevancy map. This relevancy map can support downstream tasks, such as semantic scene completion and object localization (ha2022semantic). Since CLIP can only provide image-level features, the relevancy map lacks precise pixel-level object boundary information. 3DOVS (liu2023weakly) incorporates DINO features into the NeRF output to distill object boundary information. OV-NeRF (liao2024ov) addresses the issues of coarse relevancy maps and view-inconsistent relevancy maps through SAM and cross-view self-enhancement. FMGS (zuo2025fmgs) transfers this concept from NeRF to 3DGS, achieving 851X faster inference. Although foundation models, such as CLIP and DINO, enable 3D open-set semantic scene understanding, the performance is limited by the foundation models themselves. For example, CLIP is constrained by the bag-of-words limitation (kerr2023lerf).
@@
The image features output by NeRF can be used to build a relevancy map. They can also be lifted into 3D space through multi-view images, serving as 3D features for downstream tasks (ze2023gnfactor). 3D-LLM (hong20233d) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (jatavallabhula2023conceptfusion), or Neural Field methods (hong20233d), endowing 3D features with semantic information.
@@
For implicit 3D editing, some methods use human scribbles to edit 3D shape and appearance (zhang2023nerflets; schwarz2020graf; li2024interactive; liu2021editing). However, this approach is not intuitive enough. With the development of foundation models, many methods for implicit 3D editing using text prompts have emerged. CLIP-NeRF (wang2022clip) integrates semantic features extracted by CLIP into NeRF reconstruction to change object shape and appearance during rendering. However, CLIP-based approaches cannot precisely modify specific local regions. Instruct-NeRF2NeRF (haque2023instruct) utilizes InstructPix2Pix (brooks2023instructpix2pix) to iteratively edit multiview input images and optimize the underlying scene in NeRF. This process produces a refined 3D scene that adheres to the edit instruction. However, InstructPix2Pix modifies the entire image. As a result, regions that are not desired may also be altered. DreamEditor (zhuang2023dreameditor) uses Dreambooth (ruiz2023dreambooth) to generate 2D editing masks. These masks are then converted into 3D editing regions through back projection. This approach enables precise local editing.
@@ 6.2 Pose Estimation
Object pose estimation can be divided into marker-based and markerless methods (karashchuk2021anipose). Marker-based methods require attaching passive or active markers (cassinis2007amirolos) to the object. These methods achieve high accuracy in pose estimation. For example, the NDI Polaris Vega XT, commonly used in medical robotics, can achieve an accuracy of 0.12 mm RMS (NDI_Polaris_Vega_XT). However, in unstructured environments, it is not feasible to attach specific markers to every object. Therefore, achieving object pose estimation in unstructured markerless environment is necessary.
@@
From the perspective of generalization, pose estimation methods can be classified into instance-level, category-level, and unseen object approaches (liu2024deep). Instance-level methods can estimate pose accurately for specific object instances on which they are trained. However, they struggle with novel objects. To improve the model’s adaptation for pose estimation of novel objects, category-level approaches use geometric priors from objects of the same category to estimate the pose of the novel object without requiring its 3D model (wang2019normalized). Unseen object approaches typically rely on the 3D model of the novel object to estimate its pose (caraffa2024freeze).
@@
Category-level and unseen object approaches can also be primarily classified into model-free and model-based approaches (liu2024deep). Model-free methods do not require prior knowledge of the object’s 3D model. These methods typically regress the object pose using neural network (guan2024survey). However, these methods require large amounts of data for the neural network to learn the geometric priors of the object. In contrast, model-based methods use a known 3D model of the object and usually lead the BOP benchmark for object pose estimation (burde2024comparative). However, obtaining accurate object 3D models in the real world is not easy. The advancement of multi-view image 3D reconstruction technology bridges the gap between model-based and model-free real-world applications (burde2024comparative).
@@
The input modalities for the model-based approach include RGB, depth, and RGBD. Currently, the RGBD modality leads the BOP benchmark for object pose estimation. The optimization goals are primarily divided into three parts: 2D-2D correspondences followed by regression (nguyen2024gigapose), 2D-3D correspondences followed by PnP (li2023nerf; ausserlechner2024zs6d), and 3D-3D correspondences followed by least squares fitting (lin2024sam; caraffa2024freeze). However, pose estimation accuracy remains a challenge when dealing with occlusion, specularity, symmetry and textureless objects (guan2024survey). Many methods use the predicted pose as a coarse result and refine it to obtain a fine result (labbe2022megapose; wen2024foundationpose; moon2024genflow).
@@
Figure 6: Foundation Models for State. The foundation models have three main applications in 3D reconstruction: 3D open-set semantic scene understanding, lifting 2D features to 3D space, and implicit 3D editing. In 3D open-set semantic scene understanding, the main pipeline is to use image features extracted by the VFM encoder and VLM image encoder as input for NeRF. Then, semantic text features extracted by the VLM language encoder are used in conjunction with the image features from NeRF to generate a relevancy map through a relevancy extractor (kerr2023lerf). This relevancy map can support downstream tasks, such as semantic scene completion and object localization (ha2022semantic). As for lifting, using the VLM image encoder to extract features from 3D data multi-view images and lift them into 3D features can incorporate semantic information into the 3D features. The lifting methods include direct reconstruction, gradSLAM, and Neural Field (hong20233d). For implicit 3D editing, the current mainstream pipeline is to input the image rendered by NeRF and the editing prompt into the VGM to generate the updated image. The updated image is then fed back into NeRF for training, modifying NeRF’s radiance field representation of the 3D scene (haque2023instruct). Pose estimation with foundation models achieves state-of-the-art results (caraffa2024freeze). The main method is 2D-lifting-3D. It extracts texture features from the object model and observation RGBD image. LMM extracts geometric features from the object model and observation RGBD image. The fused features are then used to estimate the 6D pose through 3D-3D correspondences.
@@
The pose estimation of moving objects mainly involves two methods. 1) Some single-image 6D pose estimation methods are fast and re-estimate poses from scratch for each frame. However, this approach is inefficient and results in less coherent estimations (wen2021bundletrack). 2) Pose tracking utilizes temporal cues to improve pose prediction. It enhances efficiency, smoothness, and accuracy in video sequences. Current pose-tracking methods are mainly divided into probabilistic tracking (deng2021poserbpf; issac2016depth; stoiber2022iterative) and optimization-based tracking (li2018deepim; wen2020se; lin2022keypoint; wang20206). Pose tracking faces challenges mainly from motion blur, incremental error drift, and occlusion. To address these issues, BundleSDF (wen2023bundlesdf) and BundleTrack (wen2021bundletrack) use an online pose graph optimization process.
@@
There are some research integrating foundation models into pose estimation. As for category-level, OV9D (cai2024ov9d) utilizes DINO and VQVAE to extract visual features from images, while CLIP is used to extract text features from category prompts. These features are then fed into the Stable Diffusion UNet (rombach2022high) to generate a normalized object coordinate space (NOCS) map. This method achieves generalizability to unseen categories and enables open-set pose estimation. In unseen object pose estimation with foundation models, FoundationPose (wen2024foundationpose) utilizes LLM-aided synthetic data generation at scale to ensure strong generalizability for novel object pose estimation & tracking. SAM-6D (lin2024sam) and ZS6D (ausserlechner2024zs6d) leverage SAM to generate valid proposals, enabling zero-shot 6D pose estimation. FreeZe (caraffa2024freeze) employs frozen GeDi (poiesi2022learning) and DINO (caron2021emerging) to extract both geometric and visual features from the query object model and the target object’s RGBD observation image. It then uses 3D-3D fused feature correspondences to obtain the 6D pose. Due to the foundation models’ robust capability in discriminative feature extraction, FreeZe achieves state-of-the-art results without the need for any data or training. Overall, foundation models primarily improve generalization for novel object pose estimation in three areas: data , object recognition, and feature extraction. However, the performance is limited by the foundation models themselves. For example, foundation models are large in size (caraffa2024freeze) and SAM may hallucinate in object segmentation (kirillov2023segment).
@@
Following Fig. 6, VLM and VFM assist implicit 3D reconstruction by generating relevancy maps that include semantic information (kerr2023lerf). They can also be employed in 2D-to-3D lifting to extract 3D features, encompassing texture, semantic, and spatial information (hong20233d). VGM aids in generating edited images and modifying 3D scenes based on these edited images (haque2023instruct). FreeZe (caraffa2024freeze) achieves state-of-the-art result in pose estimation by extracting discriminative features through 2D-to-3D lifting and LMM.